# EDNet: A Distortion-Agnostic Speech Enhancement Framework with Gating Mamba Mechanism and Phase Shift-Invariant Training

### v1

@@ III-C2 Extended Shift Search
Although the phase shift estimation method provides valid results within the limited range of $[-0.5,0.5]$, its effectiveness diminishes when the actual phase shift exceeds the range. To address this, we introduce an extended search strategy that evaluates multiple shifts before determining the optimal alignment.
@@
1:$\hat{X}_{p}$: predicted phase, $Y_{p}$: GT phase, $\mathcal{S}$: search grid
2:$n^{*}_{\text{final}}$: optimal shift
@@
[…]
@@ III-C3 Loss Computation
Once the predicted phase $\hat{X}_{p}$ is aligned using $n^{*}_{\text{final}}$, any magnitude-phase-based loss function can be directly applied to the aligned phase spectrogram, requiring no modifications beyond the alignment process itself. Following [10], we adopt the same loss composition, where the total loss is a weighted sum of the PESQ-based GAN discriminator loss, magnitude loss, phase loss, time-domain loss, complex spectrogram loss, and consistency loss. These losses are computed using the predicted magnitude and aligned phase spectrogram, with respective weights of 0.07, 27, 0.3, 6, 3, and 3.
@@ IV-A1 Denoising
To assess speech denoising performance, we adopt the VoiceBank+DEMAND dataset [60], a standard benchmark widely used in prior research. The clean speech samples originate from the VoiceBank corpus [61], comprising 11,572 utterances from 28 speakers for training, and 824 utterances from 2 unseen speakers for testing. Noise is synthetically added using 10 distinct noise types—8 sourced from the DEMAND database [62] and 2 artificially generated—at signal-to-noise ratios (SNRs) of 0, 5, 10, and 15 dB. The test set contains 5 previously unseen noise conditions at SNRs of 2.5, 7.5, 12.5, and 17.5 dB. All audio data is uniformly resampled to a 16 kHz sampling rate.
@@ IV-A2 Dereverberation
For evaluating dereverberation capabilities, we utilize the dataset provided by the REVERB Challenge [63]. The training portion comprises 7,861 reverberant utterances, generated by convolving clean WSJCAM0 [64] recordings with real room impulse responses (RIRs) and embedding ambient noise at an SNR of 20 dB. Reverberation scenarios cover three room sizes with 60 dB reverberation time (RT60) values of 0.3, 0.6, and 0.7 seconds, each captured at near (0.5 m) and far (2.0 m) microphone placements. The evaluation set consists of 2,176 simulated utterances and 372 real recordings, sourced from the MC-WSJ-AV corpus [65]. For consistency, we use the single-channel version of all recordings at 16 kHz.
@@ IV-A3 Bandwidth Extension
For the bandwidth extension task, we utilize the VCTK corpus [61], which comprises approximately 44 hours of high-quality speech data from 108 English speakers, originally recorded at a 48 kHz sampling rate. Our data preparation procedure follows established protocols from prior work [18]. Specifically, we allocate the first 100 speakers for training and reserve the remaining 8 for evaluation. To construct input-target pairs, we begin by downsampling the original 48 kHz recordings to 16 kHz, designating these as wideband reference signals. Narrowband inputs are then simulated by further subsampling the audio with factors of 2 and 4, limiting the bandwidth to 8 kHz and 4 kHz, respectively. These narrowband signals are subsequently upsampled back to 16 kHz using spline interpolation, serving as the model input. The corresponding wideband audio remains the ground truth target for supervision.
@@ IV-A4 Multi-distortion Enhancement
Since no standard benchmark exists for multi-distortion enhancement, we synthesize the data by introducing three types of distortions: additive noise, reverberation, and bandwidth limitation. Noise is mixed using the DEMAND database [62], with a randomly selected SNR ranging between -6 and 14 dB, following the same noise-type split as the VoiceBank+DEMAND dataset [60]. Reverberation is applied using room impulse responses generated by Pyroomacoustics, with room dimensions varying between 5 and 15 m in length and width, and 2 to 6 m in height. The RT60 is randomly selected between 0.4 and 1.0 seconds. Bandwidth limitation is applied using randomly selected low-pass filters, including Butterworth, Bessel, and Chebyshev filters, with cutoff frequencies set to 2, 4, or 8 kHz. We use the VCTK dataset [61] as a speech source, selecting the first 100 speakers for training and the last 8 for testing. Speakers p280 and p310 are excluded due to technical issues.
@@ IV-B Evaluation Metrics
In the denoising task, we use the perceptual evaluation of speech quality (PESQ), short-time objective intelligibility (STOI), and three composite metrics: CSIG, CBAK, and COVL [66]. PESQ provides an estimate of overall perceptual quality, while STOI quantifies intelligibility. The composite scores CSIG, CBAK, and COVL assess speech distortion, background noise intrusiveness, and overall perceptual quality, respectively.
In the dereverberation task, we adopt the official evaluation toolkit provided by the REVERB Challenge [63], which measuring various metrics including PESQ, cepstral distance (CD), log-likelihood ratio (LLR), and frequency-weighted segmental SNR (FWSegSNR). These metrics capture perceptual speech quality, time-frequency domain distortions, spectral divergence, and residual reverberation energy. Since real-recorded reverberant utterances do not have clean references, the signal-to-reverberation modulation energy ratio (SRMR) is used exclusively for those cases, serving as a non-intrusive proxy for dereverberation effectiveness.
For the bandwidth extension task, we utilize PESQ to assess the perceptual restoration of missing high-frequency components, ViSQOL (Virtual Speech Quality Objective Listener) [67] to reflect MOS-LQO (mean opinion score - listening quality objective), and log-spectral distance (LSD) [68] to quantify spectral distortion.
In the multi-distortion task, where multiple degradation types (noise, reverberation, and bandwidth limitation) co-occur, a broader and more comprehensive set of metrics is adopted. We report PESQ, CSIG, CBAK, COVL, and STOI to assess perceptual and intelligibility performance [66]. Additionally, ViSQOL [67] is used to gauge overall listening quality, while LSD [68] complements the evaluation by quantifying spectral accuracy. Finally, perceptual speaker similarity (PSS) is assessed by computing the cosine similarity between speaker embeddings of the enhanced and clean utterances. These embeddings are extracted using a WavLM-ECAPA model trained on the VoxSim dataset [69], a subset of the VoxCeleb1 dataset [70] annotated with human-labeled perceptual similarity scores.
@@ IV-C Implementation Details
All speech signals are sampled at 16 kHz. During training, random 2-second segments are extracted from each utterance. STFT is computed using a 400-point FFT, a 25 ms window (400 samples), and a 6.25 ms hop size (100 samples). Magnitude spectrograms are min-max normalized before input. All models are trained for 200 epochs using AdamW optimizer [71], with a batch size of 8 and learning rate of 0.0007, decayed exponentially by 0.99 per epoch. We track validation PESQ every 2,000 steps and select the best checkpoint. For a fair comparison and a focused analysis of the model architecture and learning objectives, we exclude specialized techniques from all experiments such as Perceptual Contrast Stretching (PCS) [72] that can boost objective score.
@@ V-A Single-Distortion Benchmark Comparison
DEMUCS [73] 3.07 4.31 3.40 3.63 0.95
CMGAN [29] 3.41 4.63 3.94 4.12 0.96
TridentSE [5] 3.47 4.70 3.81 4.10 0.96
SEMamba [11] 3.55 4.77 3.95 4.26 0.96
MP-SENet [10] 3.60 4.81 3.99 4.34 0.96
@@
WPE [74] 1.72 3.75 0.51 4.90 4.22 3.98
UNet [48] - 2.50 0.40 10.70 4.88 5.58
CMGAN [29] - 2.25 0.31 11.74 5.47 6.55
DCN [50] 2.94 2.00 0.23 13.33 5.27 6.48
MP-SENet [10] 2.97 1.97 0.24 14.07 5.51 6.67
@@
AFiLM [75] - 1.24 4.39 - 1.63 3.83
AECNN [21] 3.91 0.88 - 3.64 0.95 -
NVSR [23] 3.56 0.79 4.52 2.40 0.95 4.11
AP-BWE [22] - 0.69 4.71 - 0.87 4.30
MP-SENet [10] 4.28 0.66 4.72 3.78 0.81 4.42
@@
Denoising: Table I presents EDNet’s performance on the denoising task. EDNet achieves competitive results compared to state-of-the-art models, demonstrating strong performance across key metrics. Without PSIT, its performance is comparable to that of SEMamba [11], which shares a similar architectural foundation, combining a DenseNet-based encoder-decoder with a TF-Mamba backbone. The inclusion of PSIT yields consistent improvements in most of the key metrics. These gains suggest that PSIT can contribute to performance even in denoising scenarios, where temporal misalignment and phase distortion are relatively limited compared to other distortion types. Although STOI scores are saturated across top-performing models (0.96), the observed improvements in perceptual metrics indicate that PSIT helps produce more natural and intelligible outputs. Overall, EDNet demonstrates solid performance in denoising and serves as a strong baseline, though it has not yet surpassed all existing models, leaving room for further improvement.
@@
In summary, EDNet achieves state-of-the-art or competitive results across denoising, dereverberation, and bandwidth extension benchmarks. These results are obtained without task-specific modifications or hyperparameter tuning, demonstrating strong structural flexibility and adaptability. Beyond architectural design, PSIT consistently improves performance across various distortion types. The most substantial gains are observed in dereverberation, where temporal misalignment is pronounced. It also provides incremental but consistent benefits in scenarios involving different phase-related degradations, such as additive noise and bandwidth limitation.
@@
Noisy - 1.28 1.16 1.82 1.16 0.51 3.64 2.40 0.373
CMGAN[29] 2024 1.80 2.20 2.47 2.04 0.71 2.40 3.30 0.552
SEMamba [11] 2024 2.26 2.46 2.79 2.41 0.73 2.38 3.33 0.596
MP-SENet [10] 2025 2.24 2.63 2.75 2.49 0.72 2.31 3.25 0.604
EDNet w/o PSIT 2025 2.32 3.02 2.77 2.72 0.74 2.19 3.52 0.667
EDNet 2025 2.39 3.06 2.81 2.78 0.75 2.16 3.60 0.693
@@ V-B Multi-Distortion Performance Comparison
For multi-distortion performance comparison, we select baseline models architecturally and functionally similar to EDNet—namely, CMGAN [29], MP-SENet [10], and SeMamba [11]. These time-frequency domain models employ dilated DenseNet-based encoder-decoders with temporal-frequency modeling backbones (e.g., TF-Conformer, Transformer, Mamba). While not all target multi-distortion enhancement explicitly, their unified treatment of diverse distortions makes them suitable for evaluating generalization. All baselines offer official code, ensuring reproducibility.
@@
As summarized in Table IV, EDNet achieves the best performance across every evaluation metric. While detailed analyses are presented in following Sec. V-CV-D and V-E, the performance gains appear to be related to the complementary roles of the GM module and PSIT. The GM module employs a gating mechanism to dynamically blend masking and mapping strategies in a region-wise manner, allowing the network to emphasize the most beneficial features based on input characteristics. PSIT also contributes to performance gains in the multi-distortion setting, where diverse sources of phase degradation—such as misalignment, noise-induced perturbations, and loss of high-frequency phase information—coexist. The shift-tolerant nature of PSIT may be better aligned with these conditions, potentially allowing it to provide more consistent guidance during training. Notably, even in the absence of PSIT, EDNet consistently outperforms existing baselines across key metrics, highlighting the strength of its architectural design in multi-distortion settings.

### v2

@@ III-C2 Extended Shift Search
Although the phase shift estimation method provides valid results within the limited range of [0.5,0.5][-0.5,0.5], its effectiveness diminishes when the actual phase shift exceeds the range. To address this, we introduce an extended search strategy that evaluates multiple shifts before determining the optimal alignment.
@@
[…]
3:Initialize $n^{*}_{\text{final}}\leftarrow 0$, $\text{min\_loss}\leftarrow\infty$
@@
9:  if $\text{loss}<\text{min\_loss}$ then
10:   $\text{min\_loss}\leftarrow\text{loss}$
@@ III-C3 Loss Computation
Once the predicted phase $\hat{X}_{p}$ is aligned using $n^{*}_{\text{final}}$, any magnitude-phase-based loss function can be directly applied to the aligned phase spectrogram, requiring no modifications beyond the alignment process itself. Following [lu2023explicit], we adopt the same loss composition, where the total loss is a weighted sum of the PESQ-based GAN discriminator loss, magnitude loss, phase loss, time-domain loss, complex spectrogram loss, and consistency loss. These losses are computed using the predicted magnitude and aligned phase spectrogram, with respective weights of 0.07, 27, 0.3, 6, 3, and 3.
@@ IV-A1 Denoising
To assess speech denoising performance, we adopt the VoiceBank+DEMAND dataset [valentini2016investigating], a standard benchmark widely used in prior research. The clean speech samples originate from the VoiceBank corpus [veaux2017vctk], comprising 11,572 utterances from 28 speakers for training, and 824 utterances from 2 unseen speakers for testing. Noise is synthetically added using 10 distinct noise types—8 sourced from the DEMAND database [thiemann2013diverse] and 2 artificially generated—at signal-to-noise ratios (SNRs) of 0, 5, 10, and 15 dB. The test set contains 5 previously unseen noise conditions at SNRs of 2.5, 7.5, 12.5, and 17.5 dB. All audio data is uniformly resampled to a 16 kHz sampling rate.
@@ IV-A2 Dereverberation
For evaluating dereverberation capabilities, we utilize the dataset provided by the REVERB Challenge [kinoshita2016summary]. The training portion comprises 7,861 reverberant utterances, generated by convolving clean WSJCAM0 [fransen1994wsjcam0] recordings with real room impulse responses (RIRs) and embedding ambient noise at an SNR of 20 dB. Reverberation scenarios cover three room sizes with 60 dB reverberation time (RT60) values of 0.3, 0.6, and 0.7 seconds, each captured at near (0.5 m) and far (2.0 m) microphone placements. The evaluation set consists of 2,176 simulated utterances and 372 real recordings, sourced from the MC-WSJ-AV corpus [lincoln2005multi]. For consistency, we use the single-channel version of all recordings at 16 kHz.
@@ IV-A3 Bandwidth Extension
For the bandwidth extension task, we utilize the VCTK corpus [veaux2017vctk], which comprises approximately 44 hours of high-quality speech data from 108 English speakers, originally recorded at a 48 kHz sampling rate. Our data preparation procedure follows established protocols from prior work [kuleshov2017audio]. Specifically, we allocate the first 100 speakers for training and reserve the remaining 8 for evaluation. To construct input-target pairs, we begin by downsampling the original 48 kHz recordings to 16 kHz, designating these as wideband reference signals. Narrowband inputs are then simulated by further subsampling the audio with factors of 2 and 4, limiting the bandwidth to 8 kHz and 4 kHz, respectively. These narrowband signals are subsequently upsampled back to 16 kHz using spline interpolation, serving as the model input. The corresponding wideband audio remains the ground truth target for supervision.
@@ IV-A4 Multi-distortion Enhancement
Since no standard benchmark exists for multi-distortion enhancement, we synthesize the data by introducing three types of distortions: additive noise, reverberation, and bandwidth limitation. Noise is mixed using the DEMAND database [thiemann2013diverse], with a randomly selected SNR ranging between -6 and 14 dB, following the same noise-type split as the VoiceBank+DEMAND dataset [valentini2016investigating]. Reverberation is applied using room impulse responses generated by Pyroomacoustics, with room dimensions varying between 5 and 15 m in length and width, and 2 to 6 m in height. The RT60 is randomly selected between 0.4 and 1.0 seconds. Bandwidth limitation is applied using randomly selected low-pass filters, including Butterworth, Bessel, and Chebyshev filters, with cutoff frequencies set to 2, 4, or 8 kHz. We use the VCTK dataset [veaux2017vctk] as a speech source, selecting the first 100 speakers for training and the last 8 for testing. Speakers p280 and p310 are excluded due to technical issues.
@@ IV-B Evaluation Metrics
In the denoising task, we use the perceptual evaluation of speech quality (PESQ), short-time objective intelligibility (STOI), and three composite metrics: CSIG, CBAK, and COVL [hu2007evaluation]. PESQ provides an estimate of overall perceptual quality, while STOI quantifies intelligibility. The composite scores CSIG, CBAK, and COVL assess speech distortion, background noise intrusiveness, and overall perceptual quality, respectively.
In the dereverberation task, we adopt the official evaluation toolkit provided by the REVERB Challenge [kinoshita2016summary], which measuring various metrics including PESQ, cepstral distance (CD), log-likelihood ratio (LLR), and frequency-weighted segmental SNR (FWSegSNR). These metrics capture perceptual speech quality, time-frequency domain distortions, spectral divergence, and residual reverberation energy. Since real-recorded reverberant utterances do not have clean references, the signal-to-reverberation modulation energy ratio (SRMR) is used exclusively for those cases, serving as a non-intrusive proxy for dereverberation effectiveness.
For the bandwidth extension task, we utilize PESQ to assess the perceptual restoration of missing high-frequency components, ViSQOL (Virtual Speech Quality Objective Listener) [chinen2020visqol] to reflect MOS-LQO (mean opinion score - listening quality objective), and log-spectral distance (LSD) [gray1976distance] to quantify spectral distortion.
In the multi-distortion task, where multiple degradation types (noise, reverberation, and bandwidth limitation) co-occur, a broader and more comprehensive set of metrics is adopted. We report PESQ, CSIG, CBAK, COVL, and STOI to assess perceptual and intelligibility performance [hu2007evaluation]. Additionally, ViSQOL [chinen2020visqol] is used to gauge overall listening quality, while LSD [gray1976distance] complements the evaluation by quantifying spectral accuracy. Finally, perceptual speaker similarity (PSS) is assessed by computing the cosine similarity between speaker embeddings of the enhanced and clean utterances. These embeddings are extracted using a WavLM-ECAPA model trained on the VoxSim dataset [ahn2024voxsim], a subset of the VoxCeleb1 dataset [nagrani2020voxceleb] annotated with human-labeled perceptual similarity scores.
@@ IV-C Implementation Details
All speech signals are sampled at 16 kHz. During training, random 2-second segments are extracted from each utterance. STFT is computed using a 400-point FFT, a 25 ms window (400 samples), and a 6.25 ms hop size (100 samples). Magnitude spectrograms are min-max normalized before input. All models are trained for 200 epochs using AdamW optimizer [loshchilov2017decoupled], with a batch size of 8 and learning rate of 0.0007, decayed exponentially by 0.99 per epoch. We track validation PESQ every 2,000 steps and select the best checkpoint. For a fair comparison and a focused analysis of the model architecture and learning objectives, we exclude specialized techniques from all experiments such as Perceptual Contrast Stretching (PCS) [chao2022perceptual] that can boost objective score.
@@ V-A Single-Distortion Benchmark Comparison
DEMUCS [defossez2020real] 3.07 4.31 3.40 3.63 0.95
CMGAN [abdulatif2024cmgan] 3.41 4.63 3.94 4.12 0.96
TridentSE [yin2022tridentse] 3.47 4.70 3.81 4.10 0.96
SEMamba [chao2024investigation] 3.55 4.77 3.95 4.26 0.96
MP-SENet [lu2023explicit] 3.60 4.81 3.99 4.34 0.96
@@
WPE [nakatani2010speech] 1.72 3.75 0.51 4.90 4.22 3.98
UNet [ernst2018speech] - 2.50 0.40 10.70 4.88 5.58
CMGAN [abdulatif2024cmgan] - 2.25 0.31 11.74 5.47 6.55
DCN [kothapally2024monaural] 2.94 2.00 0.23 13.33 5.27 6.48
MP-SENet [lu2023explicit] 2.97 1.97 0.24 14.07 5.51 6.67
@@
AFiLM [rakotonirina2021self] - 1.24 4.39 - 1.63 3.83
AECNN [wang2021towards] 3.91 0.88 - 3.64 0.95 -
NVSR [liu2022neural] 3.56 0.79 4.52 2.40 0.95 4.11
AP-BWE [lu2024towards] - 0.69 4.71 - 0.87 4.30
MP-SENet [lu2023explicit] 4.28 0.66 4.72 3.78 0.81 4.42
@@
Denoising: Table I presents EDNet’s performance on the denoising task. EDNet achieves competitive results compared to state-of-the-art models, demonstrating strong performance across key metrics. Without PSIT, its performance is comparable to that of SEMamba [chao2024investigation], which shares a similar architectural foundation, combining a DenseNet-based encoder-decoder with a TF-Mamba backbone. However, it still underperforms MP-SENet [lu2023explicit], as the GM module applies masking at the feature level rather than directly in the STFT domain, which likely limits its ability to preserve fine-grained signal details. The inclusion of PSIT yields consistent improvements in most of the key metrics. These gains suggest that PSIT can contribute to performance even in denoising scenarios, where temporal misalignment and phase distortion are relatively limited compared to other distortion types. Although STOI scores are saturated across top-performing models (0.96), the observed improvements in perceptual metrics indicate that PSIT helps produce more natural and intelligible outputs. Overall, EDNet demonstrates solid performance in denoising and serves as a strong baseline, though it has not yet surpassed all existing models, leaving room for further improvement.
@@
In summary, EDNet achieves state-of-the-art or competitive results across denoising, dereverberation, and bandwidth extension benchmarks. These results are obtained without task-specific modifications or hyperparameter tuning, demonstrating strong structural flexibility and adaptability. This adaptability is most pronounced in the dereverberation task. While mapping-based reconstruction is commonly favored for dereverberation given its convolutive distortion model [lemercier2023extending, kothapally2024monaural], a single method may not be uniformly optimal across all conditions. Depending on the reverberation level, speech content may remain partially intact, favoring suppression-based strategies, whereas severe smearing can necessitate reconstruction. The strength of EDNet lies in the architectural capacity to adaptively balance these strategies rather than rely on a fixed processing paradigm. Strong results against competitive baselines from both methodological families, including the mapping-based DCN and the masking-oriented MP-SENet, support the effectiveness of this data-driven hybrid design. Beyond architectural design, PSIT consistently improves performance across various distortion types. The most substantial gains are observed in dereverberation, where temporal misalignment is pronounced. It also provides incremental but consistent benefits in scenarios involving different phase-related degradations, such as additive noise and bandwidth limitation.
@@
Noisy - - 1.28 1.16 1.82 1.16 0.51 3.64 2.40 0.373
VoiceFixer[liu2022voicefixer] 2022 122M 1.39 2.24 2.00 1.86 0.61 2.48 3.11 0.604
UNIVERSE[serra2022universal] 2022 107M 1.38 2.11 2.16 1.78 0.54 2.73 2.94 0.611
HD-DEMUCS[kim2023hd] 2023 23.6M 1.38 2.69 2.20 2.08 0.64 2.22 3.28 0.575
UNIVERSE++[scheibler2024universal] 2024 107M 1.50 2.24 2.27 1.91 0.61 2.74 3.13 0.678
CMGAN[abdulatif2024cmgan] 2024 1.83M 1.80 2.20 2.47 2.04 0.71 2.40 3.30 0.552
SEMamba [chao2024investigation] 2024 2.25M 2.26 2.46 2.79 2.41 0.73 2.38 3.33 0.596
MP-SENet [lu2023explicit] 2025 2.26M 2.24 2.63 2.75 2.49 0.72 2.31 3.25 0.604
EDNet w/o PSIT 2025 2.67M 2.32 3.02 2.77 2.72 0.74 2.19 3.52 0.667
EDNet 2025 2.67M 2.39 3.06 2.81 2.78 0.75 2.16 3.60 0.693
TABLE V: Results of Subjective Listening Test for Multi-distortion Enhancement. SQ: Speech Quality, CP: Content Preservation
[…]
VoiceFixer[liu2022voicefixer] G 1.39 2.897 3.047 2.26 $\pm$ 0.09 2.36 $\pm$ 0.07
UNIVERSE[serra2022universal] G 1.38 2.821 4.081 2.40 $\pm$ 0.09 1.83 $\pm$ 0.04
UNIVERSE++[scheibler2024universal] G 1.50 2.897 4.216 2.70 $\pm$ 0.08 2.29 $\pm$ 0.06
HD-DEMUCS[kim2023hd] D 1.38 2.712 1.240 1.41 $\pm$ 0.05 2.04 $\pm$ 0.05
CMGAN[abdulatif2024cmgan] D 1.80 2.968 2.507 2.07 $\pm$ 0.08 3.00 $\pm$ 0.06
SEMamba [chao2024investigation] D 2.26 2.968 2.885 2.56 $\pm$ 0.07 3.53 $\pm$ 0.06
MP-SENet [lu2023explicit] D 2.24 2.953 2.804 2.58 $\pm$ 0.08 3.53 $\pm$ 0.06
EDNet D 2.39 2.981 2.967 2.73 $\pm$ 0.07 3.58 $\pm$ 0.06
@@ V-B Multi-Distortion Performance Comparison
For multi-distortion performance comparison, we first consider representative multi-distortion speech enhancement models that explicitly aim to handle diverse distortion types—VoiceFixer [liu2022voicefixer], UNIVERSE [serra2022universal], HD-DEMUCS [kim2023hd], and UNIVERSE++ [scheibler2024universal]. In addition to these multi-distortion enhancement baselines, we also include CMGAN [abdulatif2024cmgan], SEMamba [chao2024investigation], and MP-SENet [lu2023explicit], which are architecturally and functionally similar to EDNet. These time–frequency domain models employ dilated DenseNet-based encoder–decoders with temporal–frequency modeling backbones such as TF-Conformer, Transformer, or Mamba. While not all target multi-distortion enhancement explicitly, their unified treatment of diverse distortions makes them suitable for evaluating generalization. All baselines offer publicly available implementations, ensuring reproducible benchmarking under consistent settings. Since VoiceFixer is designed for 44.1 kHz full-band audio, retraining it at 16 kHz is not straightforward due to architectural and preprocessing constraints. Therefore, we use its official checkpoint for inference and downsample the generated 44.1 kHz outputs to 16 kHz for evaluation. Although this setup may not represent a perfectly aligned condition, it still provides a meaningful reference for comparison, as the released checkpoint is trained with identical speech and noise sources.
@@
As summarized in Table IV, EDNet achieves the best performance across every evaluation metric. While detailed analyses are presented in following Sec. V-CV-D and V-F, the performance gains appear to be related to the complementary roles of the GM module and PSIT. The GM module employs a gating mechanism to dynamically blend masking and mapping strategies in a region-wise manner, allowing the network to emphasize the most beneficial features based on input characteristics. PSIT also contributes to performance gains in the multi-distortion setting, where diverse sources of phase degradation—such as misalignment, noise-induced perturbations, and loss of high-frequency phase information—coexist. The shift-tolerant nature of PSIT may be better aligned with these conditions, potentially allowing it to provide more consistent guidance during training. Notably, even in the absence of PSIT, EDNet consistently outperforms existing baselines across key metrics, highlighting the strength of its architectural design in multi-distortion settings.
In addition to objective evaluations, we also conduct subjective listening tests to complement the quantitative results and to better understand the perceptual differences that may not be captured by conventional metrics. Recent studies [liu2022neural, zhang2024urgent, richter2023speech] report that generative models often produce perceptually high-quality outputs despite achieving lower scores on reference-based objective metrics such as PESQ. To further analyze this discrepancy, we design two complementary listening tests: (1) speech quality test, where participants rate the perceptual quality of enhanced signals without reference access, and (2) content preservation test, where they compare the enhanced and ground-truth speechs to assess the degree of content consistency, including speaker similarity and linguistic alignment. Each test involve 20 participants and 15 randomly selected utterances from the test set. To further align these subjective assessments with objective measures, we report DNSMOS [reddy2021dnsmos] and NISQA [mittag2021nisqa] as reference-free metrics.
As summarized in Table V, generative models generally receive speech quality scores that are higher than expected from their reference-based objective metrics, highlighting a clear discrepancy between perceptual and quantitative evaluations. This suggests that their outputs often sound more natural and artifact-free than the reference-based objective scores alone indicate, a trend confirmed by their superior NISQA scores compared to all discriminative baselines. However, the same models—particularly diffusion-based systems such as UNIVERSE and UNIVERSE++—exhibit substantially lower content preservation scores than discriminative baselines, indicating frequent modifications of linguistic content or speaker identity, which likely explains their relatively low reference-based metric results. By contrast, discriminative models preserve content more faithfully but are observed to introduce high-frequency artifacts when restoring severely degraded regions, which is reflected in their relatively lower NISQA performance. Taken together, these results highlight an inherent trade-off between perceptual naturalness and content fidelity in current speech enhancement paradigms and underscore the limitations of relying solely on objective metrics for direct comparisons between discriminative and generative approaches.
EDNet achieves the best scores on both speech quality and content preservation, indicating high perceptual quality with strong content fidelity. In terms of reference-free metrics, it achieves the highest DNSMOS score overall and the highest NISQA score among the discriminative models. While its NISQA performance is still lower than that of generative models—likely due to high-frequency artifacts that are not fully resolved—EDNet shows a clear improvement over other discriminative baselines. These results suggest that our architectural design, which combines dynamic gated fusion with phase-tolerant supervision, successfully enhances perceptual quality while maintaining the reliable fidelity inherent in discriminative methods. Taken together, these results suggest a practical path to unifying the perceptual strengths of generative models with the content fidelity of discriminative methods. We further provide audio samples generated by our method on the demo page [^1] https://mm.kaist.ac.kr/projects/EDNet.